OpenStudy (anonymous):

Pls. Help I have an exam today In an arithmetic progression in which the 6th term is 4 times the 1st term. Sum of 4th term and 7th term is 37. Find the sum of all the terms from 25th to 50th.

OpenStudy (mathmale):

Might be worth reviewing just what an "arithmetic progression" is. If you wish, see: https://www.google.com/search?sourceid=chrome-psyapi2&ion=1&espv=2&es_th=1&ie=UTF-8&q=arithmetic%20progression&oq=arithmetic%20progression&aqs=chrome..69i57j0l5.3935j0j7 One approach would be to let the first term be a and the spacing between terms be d. First term: a Second: a+d Third: a+2d and so on Using these representations and the given info may help you solve the problem at hand.

OpenStudy (anonymous):

I know what an arithmetic progression is and I know these 2 equations S=n/2(a+l) S=n/2(2a+(n-1)d)

OpenStudy (anonymous):

But I still can't find it

OpenStudy (anonymous):

Because it says 25th to 50th terms